OpenStudy (anonymous):

find the equation of the line (-1,-5) and is parallel tothe line 7x+3y=-2

OpenStudy (anonymous):

y = -7/3 x -2/3 slope i s -7/3 y = -7/3 x + b -5 = 7/3 + b b = -22/3 y = -7/3 x + -22/3